OpenStudy (anonymous):

Find the equation of the line with slope 2/3 and y-intercept 4

OpenStudy (anonymous):

@coolaidd What is slope- intercept form of linear equation?

OpenStudy (anonymous):

y=2/3+4?

OpenStudy (anonymous):

Almost there, except you miss x!

OpenStudy (anonymous):

Without x, y becomes constant here :/

OpenStudy (anonymous):

so 4=2/3+4 ?

OpenStudy (anonymous):

Memorize it as long as you can, and as soon as possible y = mx + b m is slope and b is y intercept!

OpenStudy (anonymous):

y=2/3(4)+4

OpenStudy (anonymous):

You have only 2 numbers to plug in 2 places !!!

OpenStudy (anonymous):

y = m x + b m is slope and b is y intercept!

OpenStudy (anonymous):

2/3+4 !!

OpenStudy (anonymous):